VIBRONIC SPECTROSCOPY OF JET-COOLED 1,4-PHENYLENE DIISOCYANIDE

\underline{DEEPALI N. MEHTA}, ANNA K. GUTBERLET, and TIMOTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907.

This talk will present results of a gas phase, jet-cooled vibronic spectroscopy study of 1,4-PDI. A fluorescence excitation spectrum spanning the range 35,500-44,100 cm^{−1} (226.5-281.5 nm) has been recorded, as well as a resonant two photon ionization excitation and UV hole-burning spectrum, in the region of 35,500-36,500 cm^{−1}. The S_{0}-S_{1} origin is a weak band located at 35,566 cm^{−1}. The spectrum in the S_{1} region is dominated by vibronic coupling to the S_{2} state. Dispersed fluorescence spectra will be presented that uncover and characterize the vibronic coupling. We have also located the S_{0}-S_{2} origin 5687 cm^{−1} above the S_{0}-S_{1} origin. Evidence will be presented for the coupling of the S_{2} levels with nearby S_{1} levels and for internal conversion from S_{2} to S_{1} on the timescale of the fluorescence.